OpenStudy (anonymous):

In the metric system, divisions and multiples are based on the number _____?

Mehek (mehek14):

Do you have any options?

OpenStudy (frostbite):

It is based on the number 10. To give you an example we can think about the unit gram. By using the number 10 raised to the power of any other number we are able to derive all other units. For gram we for example get get: 1 megagram = 1 gram x \(10^6\) 1 kilogram = 1 gram x \(10^3\) 1 gram = 1 gram x \(10^0\) 1 decigram = 1 gram x \(10^{-1}\) 1 milligram = 1 gram x \(10^{-3}\) 1 microgram = 1 gram x \(10^{-6}\) 1 microgram = 1 gram x \(10^{-9}\) 1 picogram = 1 gram x \(10^{-12}\) It should however be noted that this doesn't only work for gram. You can derived any unit from any unit using the number 10 to the power of the correct number so long the unit is a measure for the same thing. Further you can see all the all units contain the word gram which is the basic unit, all the other units got a word before them and these we call multiplicative prefixes.